Please expalin this equation to me working this out and doing quadratic and square rooting confuses me, Solve: Solve: y(9+y)=4(2y+5) so far i have done this.
yx9+yxy=4x2+4xy+4x5
then I went 9y+y^2= 8+4y+20
from here I'm lost, Help!

Answer by mananth(16946)

y(9+y)=4(2y+5)
..
9y+y^2=8y+20
add -8y to both sides
9y-8y+y^2=8y-8y+20
y+y^2=20
add -20 to both sides
y+y^2-20=20-20
y+y^2-20=0
re organize the terms
y^2+y-20=0
y^2+5y-4y-20=0
y(y+5)-4(y+5)=0
(y-4)(y+5)=0
y= 4 OR y =-5
...
m.ananth@hotmail.ca